OpenStudy (anonymous):

How many arithmetic progressions of length 9 are there, such that the terms are integers from 1 to 70?

OpenStudy (blockcolder):

Since the arithmetic progression has 9 terms, let the first term be \(a\) and the last term be \(a+8d\), where \(d\) is the common difference. Thus, your question is equivalent to finding the number of non-negative integer solutions to \(a+8d=k\) where k goes from 9 to 70.

OpenStudy (blockcolder):

Correction: The answer to your question is actually double the answer of my equivalent question, since we also need to consider negative values of \(d\).

OpenStudy (anonymous):

OK, so the arithmetic progression has 9 terms, like a_1, a_2, a_3, ... , a_9 And the difference between each pair of terms is fixed, the same number Right? @thongkun

OpenStudy (anonymous):

How many arithmetic progressions of length 9 are there, such that the terms are integers from 1 to 70? Details and assumptions An Arithmetic Progression is a sequence of terms such that the difference between the consecutive terms is a constant. (1,2,3,4),(1,1,1,1),(4,3,2,1) are all valid arithmetic progressions of length 4.

OpenStudy (blockcolder):

The best I could come up with is a recursive formula. Let \(b_k\) be the number of non-negative solutions to \(a+8d=k\). We can prove that the generating function of \(b_k\) is \[B(x)=\frac{x}{1-x} \cdot \frac{1}{1-x^8}\] by considering the series expansion of \(B(x)\): \[B(x)=(x+x^2+x^3+\dots)(1+x^8+x^{16}+\dots).\] The exponent of \(x^k\) in this expansion is of the form \(c_1+8c_2\).

OpenStudy (anonymous):

Also, every term a_i must be an integer Also, a_1 = 1 and a_9 = 70 Now a_2 = a_1 + d, a_3 = a_2 +d, ... , a_9 = a_8 + d where d is some integer number Now, it follows from here that a_9 = a_1 + 8 *d Because you add d 8 times to go from a_1 to a_9 Now, d = (a_9 - a_1)/8 d = 69/8 = 8.625 and d is not an integer so, for example, a_2 = 9.625 is not an integer ! So, a_2 does not satisfy the condition of the problem It seems like there are no such progressions, so that it has 9 terms, from 1 to 70, and all of them integers

OpenStudy (anonymous):

Well, maybe they meant that a_1 > 1 and a_9 < 70 ? Not equal to 1 and 70, but all 9 terms lie within interval 1 to 70 ?

OpenStudy (blockcolder):

Silly me, I immediately assumed that it's just as @ⒶArchie☁✪ last said.

OpenStudy (anonymous):

I am not quite sure, it may be that they mean the second case, then we can try to count all possible progressions. tHe formulation is not exactly clear (the formulation of initial problem) So, if a_1 =1 and a_9 = 70, then there are no such progressions, but in the second case, surely there are some And we can count them

OpenStudy (anonymous):

For example, one will be a_1 = 1, a_2 = 2, a_3 = 3, a_4 = 4 , ... , a_9 = 9 we can count first by going through all values of d d must be integer from 1 to 8 ! Right ? d cannot be larger than 8.625 that we just obtained Because, if d > 8.625 then all 9 terms will not fit into the interval from 1 to 70 because 8 * d < 70 -1 8 * d < 69 right ?

OpenStudy (anonymous):

Ok, whatever a_1 is, it is always true that a_9 - a_1 = 8 * d Do you agree ? @blockcolder

OpenStudy (anonymous):

This is from the definition of arithmetic progression

OpenStudy (anonymous):

a_1 is the first term, a_1 + d is the second term, ..., a_9 = a_1 + 8d is the 9th term

OpenStudy (blockcolder):

Yeah, but a_1 can be anything from 1 to 70, and d can be anything from -8 to 8. Moreover, some combinations of a_1 and d are incompatible due to the restriction that the terms be all between 1 and 70. How do you plan to count them?

OpenStudy (anonymous):

How do you plan to count them? (y) :)

OpenStudy (anonymous):

@blockcolder I have this plan. Every progression is uniquely defined by a_1 and d. So, we go through all possible d and a_1 and count them all. First, we start with d's, and I thought that negative ds are not considered But if you want, we can consider negative d as well Then, we have 8 possible choices for positive d, d =1, d =2, d = 3, ... , d =8

OpenStudy (anonymous):

@thongkun

OpenStudy (blockcolder):

Well, the negative d is for sequences like 9,8,7,...,1.

OpenStudy (anonymous):

Now for every choice of d we will count all possible choices of a_1 For example, if d =1, then a_1 can be any integer from 1 to 62 Because if a_1 > 62, then a_9 will be larger than 62 + 8 = 70 So, a_9 will be larger than 70, and will not satisfy conditions of the problem But any a_1 from 1 to 62 is OK, So, we have 62 possible choices for a_1 and 61 possible sequences with d =1 Those are unique progressions, and we count them all Then, we go to d =2 ... and d =3 .. and finally d = 8 This is my plan

OpenStudy (anonymous):

I think the answer to this problem dosent acquire negative d

OpenStudy (anonymous):

@thongkun @blockcolder What do you think on this plan?

OpenStudy (anonymous):

also @blockcolder Right ... we will deal with negative d as well later

OpenStudy (anonymous):

The number of progressions with negative d is just equal to number of progressions with positive d Did you consider negative d in class ? Or only positive d, when you were discussing arithmetic progressions ? @thongkun

OpenStudy (anonymous):

wait me for understand your plan. It seems good. :)

OpenStudy (anonymous):

Even, if we add negative d, the number of progressions will be just double ... it is easy to take them into account

OpenStudy (blockcolder):

That's actually a nice idea. It's better than my generating functions at least. Here's a medal for you. :)

OpenStudy (anonymous):

Thank you! o.o

OpenStudy (anonymous):

I am glad you guys understand the plan and agree on it

OpenStudy (anonymous):

There might be other ways to solve this problem ... I just think of this plan at the moment, it is the first idea that came to my mind So, let us start counting

OpenStudy (anonymous):

I think d =0 is acceptable in the problem.

OpenStudy (anonymous):

If d =0, then all terms are the same OK, then the number of progressions with d = 0 is 70

OpenStudy (anonymous):

Because this is just the number of integers from 1 to 70

OpenStudy (anonymous):

If you want to count those single numbered progressions with d = 0 then it is just 70

OpenStudy (anonymous):